OpenStudy (anonymous):

What is the sum of a 7–term geometric series if the first term is –11, the last term is –45,056, and the common ratio is –4?

OpenStudy (anonymous):

i got -36,047?

OpenStudy (anonymous):

sum of n terms = a (r^n - 1) ----- ( r - 1) = -11 * ((- 4)^7 - 1) --------- - 4 - 1

OpenStudy (anonymous):

thank you so much for showing me how.

OpenStudy (anonymous):

can i ask your help with one more?

OpenStudy (anonymous):

i get that -36047

OpenStudy (anonymous):

i might have time so try me

OpenStudy (anonymous):

What is the sum of the geometric sequence –4, 24, –144, … if there are 7 terms?

OpenStudy (anonymous):

thank you so much!